.

A rectangular metal waveguide filled with a dielectric of relative permittivity εr = 4, has the inside dimensions 3 x 1.2 cm, the cut off frequency for the dominant mode is


1
Expert's answer
2021-01-06T13:53:15-0500

Answer

"V_p = \frac{c }{\sqrt{(μ_rε_r)} }\\here \\μ_r =1 ; ε_r = 4 (given).
\\So \\f_c =\frac{V_p}{2a}
\\f_c=\frac{3\times10^8}{((3\times10^{-2})\sqrt{(1\times4}) }\\f_c=2.5GHz"